A 90% antifreeze solution is to be mixed with a 75% solution to get 120 liters of a 78% solution. How many liters of the 90% and 75% solutions will be used?


Thank you in advance! Hoping for your immediate reply.

Answer by stanbon(75887)

A 90% antifreeze solution is to be mixed with a 75% solution to get 120 liters of a 78% solution. How many liters of the 90% and 75% solutions will be used?
-----------
Equation:
active + active = active
0.9x + 0.75(120-x) = 0.78*120
---
Multiply thru by 100 to get:
90x + 75*120 - 75x = 78*120
15x = 3*120
x = (1/5)*120
x = 24 liters (amt of 90% solution needed)
120-x = 96 liter (amt of 75% solution needed)
=========
Cheers,
Stan H.
=========